# Energy exchange between electrons and ions in ion temperature gradient turbulence

### v1

@@ III.3 Comparison between collisional and turbulent energy exchanges
Figure 3: The wavenumber spectra of the linear growth rate and frequency (a) and the turbulent energy transfer terms in Eqs. (24)–(28) for electrons (b) and ions (c) in the case of $T_{e}/T_{i}=1.0$. The spectra are given as functions of $k_{y}\rho_{ti}$ obtained by summing over $k_{x}$. The peaks and valleys of the turbulent energy transfer terms are found at a wavenumber lower than that of the linearly most unstable mode. The directions and magnitudes of the turbulent energy transfer terms are represented by arrows in the bottom figure which schematically shows the role of turbulent energy transfer terms based on Eqs. (18), (24), and (II.3) with magnetic fluctuations neglected. The electron heating due to the parallel field denoted by $Q_{e\parallel}^{\rm turb}>0$ and the ion cooling due to the $\nabla B$-curvature drift denoted by $Q_{iB}^{\rm turb}<0$ are dominant mechanisms in the turbulent energy exchange between electrons and ions in ITG turbulence.
@@
It can also be confirmed that as the plasma temperature is increased, the energy exchange by turbulence becomes more dominant than that by collision. Note that, based on the gyrokinetic ordering, the turbulent ion heating is written as $Q_{i}^{\rm turb}=(\rho_{ti}/R_{0})^{2}(v_{ti}/R_{0})(nT_{i})Q^{*}$ where $Q^{*}$ is a dimensionless function of dimensionless variables $T_{e}/T_{i},R_{0}/L_{n},R_{0}/L_{Ti},R_{0}/L_{Te},q_{0},\hat{s},\beta,R_{0}\nu_{i}/v_{ti},\cdots$ which are used as input parameters of the local flux-tube simulation. This expression is combined with Eq. (21) to obtain
@@ III.4 Spectral analysis of the turbulent energy exchange
Next, we examine each component of the wavenumber spectrum of the turbulent energy exchange shown in Eq.(24). Figure 3 (a) shows the wavenumber spectra of the linear growth rate and frequency. The turbulent energy transfer terms in Eqs. (24)–(28) in the case of $T_{e}/T_{i}=1.0$ are shown for electrons and ions in Figs. 3 (b) and (c), respectively. First, the collisional components $Q^{\mathrm{turb}}_{eC\bm{k}_{\perp}}$ and $Q^{\mathrm{turb}}_{iC\bm{k}_{\perp}}$ have little effect on the turbulent energy exchange. In addition, both electrons and ions show positive values of the parallel-heating components $Q^{\mathrm{turb}}_{s\parallel\bm{k}_{\perp}}>0$ $(s=e,i)$ , and negative values of the components $Q^{\mathrm{turb}}_{sB\bm{k}_{\perp}}<0$ $(s=e,i)$ caused by the product of the perpendicular field and the $\nabla B$curvature drift velocity in the wavenumber region where the ITG mode is linearly unstable. At the bottom of Fig.3, the roles of turbulent energy transfer terms in increasing or decreasing the thermal electron and ion energies, and the turbulent electrostatic electric field energy are schematically shown, based on Eqs. (18), (24), and (II.3) with the magnetic fluctuations neglected. We can see that the perpendicular ion cooling represented by $Q^{\mathrm{turb}}_{iB}<0$ is dominant and overcomes the parallel ion heating $Q^{\mathrm{turb}}_{i\parallel}>0$, which leads to the net turbulent ion cooling shown by $Q_{i}^{\rm turb}<0$. On the other hand, the parallel heating $Q^{\mathrm{turb}}_{e\parallel}>0$ is dominant for electrons. Thus, the perpendicular ion cooling and the parallel electron heating are found to be the main mechanisms in the turbulent energy transfer from ions to electrons in a steady state of ITG turbulence.
@@
This perpendicular ion cooling is connected with the mechanism of ITG instability. In low beta plasmas, the $\nabla B$-curvature drift can be expressed as $\bm{v}_{sd}=\bm{b}\times\left(v_{\parallel}^{2}+\mu B/m_{s}\right)\nabla B/\Omega_{s}B$. Then, under the electrostatic approximation, $Q^{\mathrm{turb}}_{iB\bm{k}_{\perp}}$ in Eq. (26) is rewritten as
@@
$Q^{\mathrm{turb}}_{iB\bm{k}_{\perp}}=\operatorname{\rm{Re}}\Bigg{\langle}\Bigg{\langle}2P_{i\bm{k}_{\perp}}^{*}\left(\frac{c}{B}\bm{E}_{\bm{k}_{\perp}}\times\bm{b}\right)\cdot\nabla B\Bigg{\rangle}\Bigg{\rangle}$ (50)
@@
where $\bm{E}_{\bm{k}_{\perp}}\equiv-i\bm{k}_{\perp}\hat{\phi}$ is the electric field and $P_{i\bm{k}_{\perp}}\equiv\int d^{3}v\left(m_{i}v_{\parallel}^{2}+\mu B\right)f_{i\bm{k}_{\perp}}/2$ roughly represents ion pressure perturbation. The ITG mode is destabilized at the outside of the torus (or the bad curvature region) where the ion pressure perturbation is amplified by the outward ${\bf E}\times{\bf B}$ flow from the inner hot plasma region. Therefore, the phases of $P_{i\bm{k}_{\perp}}$ and $(c\bm{E}_{\bm{k}_{\perp}}\times\bm{b}/B)\cdot\nabla B$ become opposite to each other, and accordingly, $Q^{\mathrm{turb}}_{iB}<0$ is expected for the ITG instability from Eq. (50). We also see from Eq. (50) that $-Q^{\mathrm{turb}}_{iB}$ represents the outward energy flow in the bad curvature. Thus, $Q^{\mathrm{turb}}_{iB}<0$ means the outward energy transport $-Q^{\mathrm{turb}}_{iB}>0$ due to the ITG turbulence.
@@
In Fig. 3, the nonlinear interaction between different wavenumbers is shown by the components $Q^{\mathrm{turb}}_{e\psi\bm{k}_{\perp}}$ and $Q^{\mathrm{turb}}_{i\psi\bm{k}_{\perp}}$. They show that the energy in the unstable wavenumber region is carried to the zonal flow mode with $k_{y}=0$. This implies that the zonal flows play a significant role in the nonlinear saturation of the ITG turbulence. It is also confirmed that $\sum_{\bm{k}_{\perp}}Q^{\mathrm{turb}}_{s\psi\bm{k}_{\perp}}=0$ $(s=e,i)$ and the nonlinear interaction causes no net energy production.
@@ IV Conclusions and Discussion
Since the effect of turbulence on the energy exchange between electrons and ions can possibly overcome that of Coulomb collisions, the turbulent energy exchange as well as the turbulent transport fluxes of particles and heat should be taken into account for predicting global profiles of the density and temperature in future fusion reactors. In order to examine the predictability of the quasilinear model of the energy exchange and transport fluxes, the quasilinear weights of these turbulent quantities normalized by the squared potential are estimated as functions of the wavenumber by linear simulations. It is found that the quasilinear weights of both energy exchange and fluxes agree with the nonlinear simulation results within an error margin of $30\%$ in the wavenumber region, where more than $80\%$ of total energy exchange and fluxes are covered. This indicates that we can construct the quasilinear model which is valid for predicting the energy exchange as well as the transport fluxes.
@@ acknowledgements
The present study is supported in part by the JSPS Grants-in-Aid for Scientific Research Grant No. 19H01879 and in part by the NIFS Collaborative Research Program NIFS23KIPT009. This work is also supported by JST SPRING, Grant Number JPMJSP2108. Simulations in this work were performed on “Plasma Simulator” (NEC SX-Aurora TSUBASA) of NIFS with the support and under the auspices of the NIFS Collaboration Research program (NIFS23KIPT009).

### v2

@@ III.3 Comparison between collisional and turbulent energy exchanges
Figure 3: The wavenumber spectra of the linear growth rate and frequency (a) and the turbulent energy transfer terms in Eqs. (24)–(28) for electrons (b) and ions (c) in the case of $T_{e}/T_{i}=1.0$. The spectra are given as functions of $k_{y}\rho_{ti}$ obtained by summing over $k_{x}$. The peaks and valleys of the turbulent energy transfer terms are found at a wavenumber lower than that of the linearly most unstable mode. The directions and magnitudes of the turbulent energy transfer terms are represented by arrows in the bottom figure which schematically shows the role of turbulent energy transfer terms based on Eqs. (II.2), (24), and (II.3) with magnetic fluctuations neglected. The electron heating due to the parallel field denoted by $Q_{e\parallel}^{\rm turb}>0$ and the ion cooling due to the $\nabla B$-curvature drift denoted by $Q_{iB}^{\rm turb}<0$ are dominant mechanisms in the turbulent energy exchange between electrons and ions in ITG turbulence.
@@
It can also be confirmed by comparing with Figs. 2 (a) and (b) that as the temperatures for electrons and ions are increased with their ratio $T_{e}/T_{i}$ fixed, the energy exchange by turbulence becomes more dominant than that by collision. Note that, based on the gyrokinetic ordering, the turbulent ion heating is written as $Q_{i}^{\rm turb}=(\rho_{ti}/R_{0})^{2}(v_{ti}/R_{0})(nT_{i})Q^{*}$ where $Q^{*}$ is a dimensionless function of dimensionless variables $T_{e}/T_{i},R_{0}/L_{n},R_{0}/L_{Ti},R_{0}/L_{Te},q_{0},\hat{s},\beta,R_{0}\nu_{i}/v_{ti},\cdots$ which are used as input parameters of the local flux-tube simulation. This expression is combined with Eq. (21) to obtain
@@ III.4 Spectral analysis of the turbulent energy exchange
Next, we examine each component of the wavenumber spectrum of the turbulent energy exchange shown in Eq.(24). Figure 3 (a) shows the wavenumber spectra of the linear growth rate and frequency. The turbulent energy transfer terms in Eqs. (24)–(28) in the case of $T_{e}/T_{i}=1.0$ are shown for electrons and ions in Figs. 3 (b) and (c), respectively. First, the collisional components $Q^{\mathrm{turb}}_{eC\bm{k}_{\perp}}$ and $Q^{\mathrm{turb}}_{iC\bm{k}_{\perp}}$ have little effect on the turbulent energy exchange. In addition, both electrons and ions show positive values of the parallel-heating components $Q^{\mathrm{turb}}_{a\parallel\bm{k}_{\perp}}>0$ $(a=e,i)$ , and negative values of the components $Q^{\mathrm{turb}}_{aB\bm{k}_{\perp}}<0$ $(a=e,i)$ caused by the product of the perpendicular field and the $\nabla B$curvature drift velocity in the wavenumber region where the ITG mode is linearly unstable. At the bottom of Fig.3, the roles of turbulent energy transfer terms in increasing or decreasing the thermal electron and ion energies, and the turbulent electrostatic electric field energy are schematically shown, based on Eqs. (II.2), (24), and (II.3) with the magnetic fluctuations neglected. We can see that the perpendicular ion cooling represented by $Q^{\mathrm{turb}}_{iB}<0$ is dominant and overcomes the parallel ion heating $Q^{\mathrm{turb}}_{i\parallel}>0$, which leads to the net turbulent ion cooling shown by $Q_{i}^{\rm turb}<0$. On the other hand, the parallel heating $Q^{\mathrm{turb}}_{e\parallel}>0$ is dominant for electrons. Thus, the perpendicular ion cooling and the parallel electron heating are found to be the main mechanisms in the turbulent energy transfer from ions to electrons in a steady state of ITG turbulence.
@@
This perpendicular ion cooling is connected with the mechanism of ITG instability. In low beta plasmas, the $\nabla B$-curvature drift can be expressed as $\bm{v}_{da}=\bm{b}\times\left(v_{\parallel}^{2}+\mu B/m_{a}\right)\nabla B/\Omega_{a}B$. Then, under the electrostatic approximation, $Q^{\mathrm{turb}}_{iB\bm{k}_{\perp}}$ in Eq. (26) is rewritten as
@@
$Q^{\mathrm{turb}}_{iB\bm{k}_{\perp}}=\operatorname{\rm{Re}}\Bigg{\langle}\Bigg{\langle}2P_{i\bm{k}_{\perp}}^{*}\left(\frac{c}{B}\bm{E}_{\bm{k}_{\perp}}\times\bm{b}\right)\cdot\nabla\ln{B}\Bigg{\rangle}\Bigg{\rangle}$ (50)
@@
where $\bm{E}_{\bm{k}_{\perp}}\equiv-i\bm{k}_{\perp}\hat{\phi}$ is the electric field and $P_{i\bm{k}_{\perp}}\equiv\int d^{3}v\left(m_{i}v_{\parallel}^{2}+\mu B\right)f_{i\bm{k}_{\perp}}/2$ roughly represents ion pressure perturbation. The ITG mode is destabilized at the outside of the torus (or the bad curvature region) where the ion pressure perturbation is amplified by the outward ${\bf E}\times{\bf B}$ flow from the inner hot plasma region. Therefore, the phases of $P_{i\bm{k}_{\perp}}$ and $(c\bm{E}_{\bm{k}_{\perp}}\times\bm{b}/B)\cdot\nabla\ln{B}$ become opposite to each other, and accordingly, $Q^{\mathrm{turb}}_{iB}<0$ is expected for the ITG instability from Eq. (50). We also see from Eq. (50) that $-Q^{\mathrm{turb}}_{iB}$ represents the outward energy flow in the bad curvature. Thus, $Q^{\mathrm{turb}}_{iB}<0$ means the outward energy transport $-Q^{\mathrm{turb}}_{iB}>0$ due to the ITG turbulence.
@@
In Fig. 3, the nonlinear interaction between different wavenumbers is shown by the components $Q^{\mathrm{turb}}_{e\psi\bm{k}_{\perp}}$ and $Q^{\mathrm{turb}}_{i\psi\bm{k}_{\perp}}$. They show that the energy in the unstable wavenumber region is carried to the zonal flow mode with $k_{y}=0$. This implies that the zonal flows play a significant role in the nonlinear saturation of the ITG turbulence. It is also confirmed that $\sum_{\bm{k}_{\perp}}Q^{\mathrm{turb}}_{a\psi\bm{k}_{\perp}}=0$ $(a=e,i)$ and the nonlinear interaction causes no net energy production.
@@ IV Conclusions and Discussion
Since the effect of turbulence on the energy exchange between electrons and ions can possibly overcome that of Coulomb collisions, the turbulent energy exchange as well as the turbulent transport fluxes of particles and heat should be taken into account for predicting global profiles of the density and temperature in future fusion reactors. In order to examine the predictability of the quasilinear model of the energy exchange and transport fluxes, the quasilinear weights of these turbulent quantities normalized by the squared potential are estimated as functions of the wavenumber by linear simulations. It is found that the quasilinear weights of both energy exchange and fluxes agree with the nonlinear simulation results within an error margin of $30\%$ in the wavenumber region, where more than $80\%$ of total energy exchange and fluxes are covered. This indicates that we can construct the quasilinear model which is valid for predicting the energy exchange as well as the transport fluxes. Although this study has not investigated the saturation model, it would not be difficult to incorporate turbulent energy exchange in existing codes that predict fluxes with a quasilinear modelStaebler ; Parker .
@@
Furthermore, in principle, the theory and simulation methods for turbulent energy exchange in this study are expected to be applicable or extendable to non-tokamak systems, such as helical systems (stellarators and heliotrons)Wakatani . The study of turbulent energy exchange and quasilinear weights in such general 3D systems is also the subject of future research.
@@ acknowledgements
The present study is supported in part by the JSPS Grants-in-Aid for Scientific Research Grant No. 19H01879, 24K07000, and in part by the NIFS Collaborative Research Program NIFS23KIPT009. This work is also supported by JST SPRING, Grant Number JPMJSP2108 and the NINS program of Promoting Research by Networking among Institutions, Grant Number 01422301. Simulations in this work were performed on “Plasma Simulator” (NEC SX-Aurora TSUBASA) of NIFS with the support and under the auspices of the NIFS Collaboration Research program (NIFS23KIPT009).
@@
Appendix A Discussion of the effect of electric field
In this work, we follow the local gyrokinetic theory based on the WKB (or ballooning) representationAntonsen , and assume that background $E\times B$ velocity is on the order of $(\rho_{ta}/L)v_{ta}=\delta v_{ta}$ (the same order as that of the diamagnetic velocity $\bm{v}_{*a}$) and that the gradient scale length of the background electric field is the same as the equilibrium scale length $L$. Under these assumptions, the Doppler frequency shift due to the background $E\times B$ velocity $\bm{k}_{\perp}\cdot\bm{v}_{E}$ is considered to be on the same order as the diamagnetic frequency $\bm{k}_{\perp}\cdot\bm{v}_{*a}$ although the background $E\times B$ velocity shear effect is estimated by $\Delta r(\partial v_{E}/\partial r)(\nabla r\times\bm{b})\cdot\nabla\hat{f}_{a}/|\nabla r|\sim\rho_{ta}(\delta v_{ta}/L)k_{\perp}\hat{f}_{a}\sim(\delta v_{ta}/L)\hat{f}_{a}\sim(\delta k_{\perp}v_{*a})\hat{f}_{a}$, the magnitude of which is smaller than terms included in the gyrokinetic equation, Eq. (II.1), by a factor of $\delta$. Here, $\Delta r$ is the radial correlation length of the fluctuation which is estimated as the typical perpendicular wavelength $1/k_{\perp}\sim\rho_{ta}$.
If the background $E\times B$ flow is larger and/or the gradient scale length of the background electric field is smaller than assumed above, the background $E\times B$ flow shear effect cannot be neglected and both neoclassical and turbulent processes of transport and energy exchange are considered to have $E_{r}$-dependence different from the results in this paper. Also, the coupling of neoclassical and turbulent processes which influences the $E_{r}$-dependence may occur in a global model including the sheared $E_{r}$ profile. A nonlinear gyrokinetic equation with large flow velocities on the order of the ion thermal speed is derived in Ref. Sugama1998 .
Based on the ordering assumption described above, the sum of the last two terms on the right-hand side of Eq.(II.2) can be rewritten asSugama1996
$-e_{a}\Gamma_{a}^{\rm{turb}}\frac{\partial\Phi}{\partial r}+Q^{\rm turb}_{a}$
$=e_{a}\sum_{\bm{k}_{\perp}}\operatorname{\rm{Re}}\bigg{\langle}\bigg{\langle}\int d^{3}v\bigl{[}e^{i\bm{k}_{\perp}\cdot\bm{v}_{E}}h_{a\bm{k}_{\perp}}\big{]}^{*}\frac{\partial}{\partial t}\big{[}e^{i\bm{k}_{\perp}\cdot\bm{v}_{E}}\psi_{a\bm{k}_{\perp}}\big{]}\bigg{\rangle}\bigg{\rangle}.$
As pointed out in Ref.Sugama1996 , the background radial electric field $E_{r}=-\partial\Phi/\partial r$ enters the gyrokinetic equation, Eq. (II.1) only in the form of the Doppler shift $(\partial/\partial t+i\bm{k}_{\perp}\cdot\bm{v}_{E})$ and does not appear explicitly in Eqs. (II.1)–(9). Therefore, for the solutions of Eq. (II.1) and Eqs. (7)–(9), $e^{i\bm{k}_{\perp}\cdot\bm{v}_{E}}h_{a\bm{k}_{\perp}}$ and $e^{i\bm{k}_{\perp}\cdot\bm{v}_{E}}\psi_{a\bm{k}_{\perp}}$ are independent of $E_{r}$. Then, even though each of the first and second terms on the left-hand side of Eq. (A) depends on $E_{r}$, their sum does not. In the present paper, we evaluate $Q^{\rm turb}_{a}$ by the gyrokinetic simulation for $E_{r}=0$ although we should note that the results of $Q^{\rm turb}_{a}$ for $E_{r}=0$ are equivalent to those of $e_{a}\Gamma_{a}^{\rm{turb}}E_{r}+Q^{\rm turb}_{a}$ for any $E_{r}$. We also find from Eq. (14) that the right-hand side of the entropy balance equation, Eq. (12), contains this sum of the terms $e_{a}\Gamma_{a}^{\rm{turb}}E_{r}+Q^{\rm turb}_{a}$ which takes a value independent of $E_{r}$.
Appendix B Derivation of Eq. (24)
In this Appendix , the derivation of Eq. (24) is presented. We first note
$\Bigg{\langle}\Bigg{\langle}e_{a}\int d^{3}vh_{a\bm{k}_{\perp}}^{*}\frac{\partial\psi_{a\bm{k}_{\perp}}}{\partial t}\Bigg{\rangle}\Bigg{\rangle}$ $=$ $\frac{\partial}{\partial t}\Bigg{\langle}\Bigg{\langle}e_{a}\int d^{3}v\psi_{a\bm{k}_{\perp}}h^{*}_{a\bm{k}_{\perp}}\Bigg{\rangle}\Bigg{\rangle}$
$-\Bigg{\langle}\Bigg{\langle}e_{a}\int d^{3}v\psi_{a\bm{k}_{\perp}}\frac{\partial h^{*}_{a\bm{k}_{\perp}}}{\partial t}\Bigg{\rangle}\Bigg{\rangle},$
[…]
$\Bigg{\langle}\Bigg{\langle}-e_{a}\sum_{\bm{k}_{\perp}}\int d^{3}v\psi_{a\bm{k}_{\perp}}\frac{\partial h^{*}_{a\bm{k}_{\perp}}}{\partial t}\Bigg{\rangle}\Bigg{\rangle}$
$=\Bigg{\langle}\Bigg{\langle}e_{a}\sum_{\bm{k}_{\perp}}\int d^{3}v\psi_{a\bm{k}_{\perp}}\bigg{(}v_{\parallel}\bm{b}\cdot\nabla h^{*}_{a\bm{k}_{\perp}}+i\bm{k}_{\perp}\cdot\bm{v}_{da}h^{*}_{a\bm{k}_{\perp}}$
$\hskip 10.0pt-\frac{c}{B}\sum_{\bm{k}_{\perp}^{\prime}+\bm{k}_{\perp}^{\prime\prime}=\bm{k}_{\perp}}\left[\bm{b}\cdot\left(\bm{k}_{\perp}^{\prime}\times\bm{k}_{\perp}^{\prime\prime}\right)\right]\psi^{*}_{a\bm{k}_{\perp}^{\prime}}h^{*}_{a\bm{k}_{\perp}^{\prime\prime}}-C_{a}^{GK*}\bigg{)}\Bigg{\rangle}\Bigg{\rangle}.$
On the right-hand side of Eq. (B), a pure imaginary term disappears and a time derivative term is neglected again because of the transport time scale ordering. We immediately see that Eqs.(26) and (28) corresponds to the second and fourth terms on the right-hand side of Eq.(B). Equation (27) is also derived from the third term by $\sum_{\bm{k}_{\perp}^{\prime}+\bm{k}_{\perp}^{\prime\prime}=\bm{k}_{\perp}}\left[\bm{b}\cdot\left(\bm{k}_{\perp}^{\prime}\times\bm{k}_{\perp}^{\prime\prime}\right)\right]=\sum_{\bm{k}_{\perp}^{\prime}+\bm{k}_{\perp}^{\prime\prime}=\bm{k}_{\perp}}\left[\bm{b}\cdot\left(\bm{k}_{\perp}^{\prime}\times\bm{k}_{\perp}\right)\right]$.
When phase space variables ($\bm{x}$, $w=m_{a}v^{2}/2$, $\mu=m_{a}v_{\perp}^{2}/2B$) are used, the integral over velocity space is described as
[…]
where $\xi$ is the gyrophase and $\sigma=v_{\parallel}/|v_{\parallel}|=\pm 1$. The first term can be written as
$\Bigg{\langle}\Bigg{\langle}e_{a}\int d^{3}v\psi_{a\bm{k}_{\perp}}v_{\parallel}\bm{b}\cdot\nabla h^{*}_{a\bm{k}_{\perp}}\Bigg{\rangle}\Bigg{\rangle}$
[…]
$\hskip 10.0pt-\Bigg{\langle}\Bigg{\langle}e_{a}\int d^{3}v\big{(}h^{*}_{a\bm{k}_{\perp}}v_{\parallel}\bm{b}\cdot\nabla\psi_{a\bm{k}_{\perp}}\big{)}\Bigg{\rangle}\Bigg{\rangle}$
[…]
$\hskip 10.0pt-\Bigg{\langle}\Bigg{\langle}e_{a}\int d^{3}v\big{(}h^{*}_{a\bm{k}_{\perp}}v_{\parallel}\bm{b}\cdot\nabla\psi_{a\bm{k}_{\perp}}\big{)}\Bigg{\rangle}\Bigg{\rangle}.$ (67)
Since the integral range of $\mu$ depends on position $\bm{x}$, the boundary of the integral range needs to be taken into account when moving $\nabla$ outside of the integral on the right-hand side of Eq.(B). However, when $\mu=w/B$, it is a bounce point($v_{\parallel}=0$) and the integral function $\sum_{\sigma}\sigma h^{*}_{a\bm{k}_{\perp}}\psi_{a\bm{k}_{\perp}}$ at the point is zero. Therefore, the boundary effect does not affect the integral calculations. The flux surface average of the first term of the right-hand side in Eq. (B) vanishes due to $\langle\bm{B}\cdot\nabla\cdots\rangle=0$.
Then, we obtain
$\Bigg{\langle}\Bigg{\langle}e_{a}\int d^{3}v\psi_{a\bm{k}_{\perp}}v_{\parallel}\bm{b}\cdot\nabla h^{*}_{a\bm{k}_{\perp}}\Bigg{\rangle}\Bigg{\rangle}$
$\hskip 10.0pt=\Bigg{\langle}\Bigg{\langle}-e_{a}\int d^{3}vh^{*}_{a\bm{k}_{\perp}}v_{\parallel}\bm{b}\cdot\nabla\psi_{a\bm{k}_{\perp}}\Bigg{\rangle}\Bigg{\rangle},$ (68)
which gives Eq. (25).